Please show your work and explain for me so I can understand better please. 2. The following is record of Restaurant 2313s purchases and sales of wines: March 2: Purchased 12 bottles of M&B wine @$12.
Answer #1

Under FIFO method goods purchased first are sold first. Therefore the total sales made during the month of March will be made from the purchased made first.

Therefore the cost of sales will be

12 bottles purchased on March 2 and remaining 20 bottles purchased on March 16.

= 12*12.50 + 20*13.50

= 150 + 270

= $420

Ending inventory will be from the purchase made on 20th

= 4*13.50

= $54

B)

Under LIFO method goods purchased last are sold first .

Therefore the total sales made during the month of March will be from purchase made on last.

Cost of sales will be from 24 bottles purchased on March 16 and remaining 8 bottles from purchase made on March 2

= 24*13.50 + 8*12.50

= 324 + 100

= $424

Ending inventory will be from purchase made on March 2

= 4*12.50

= $50.

If you find the answer helpful please upvote.